If x1,x2,x3 are in AP and y1,y2,y3 are also in AP,prove that the points (x1,y1),(x2,y2),(x3,y3) are collinear.

Grade:11

1 Answers

kamalesh bagrecha
7 years ago
Since x1,x2,x3 are in ap hence x2-x1=x3-x2Sumilarly Y2-y1=y3-y2Divide both eq(y2-y1)/(x2-x1)=(y3-y2)/(y2-y1)Hence points are coliear